OpenStudy (anonymous):

simplify In 9 - 4 In 3

OpenStudy (abb0t):

Apply the rules of natural log \[\ln(a) - \ln(b) = \ln(\frac{ a }{ b })\] similarly, \[p \ln (r) = \ln(r)^p \]

OpenStudy (precal):